OpenStudy (mtalhahassan2):

The population of a city is growing at an average rate of 3% per year. In 1990, the population was 45 000. a) Write an equation that models the growth of the city. Explain what each part of the equation represents. b) Use your equation to determine the population of the city in 2007. c) Determine the year during which the population will have doubled. d) Suppose the population took only 10 years to double. What growth rate would be required for this to have happened?

OpenStudy (mtalhahassan2):

Need help with part D

OpenStudy (mtalhahassan2):

@allibricker12

OpenStudy (mtalhahassan2):

@M,aliha

OpenStudy (mtalhahassan2):

@maddie456

OpenStudy (anonymous):

about 7%

OpenStudy (anonymous):

here is for all 4. a) y=45000 (1 .03)^x b) 74 378 (approx.) c) by 2013 d) about 7%

OpenStudy (mtalhahassan2):

c be 2014

OpenStudy (mtalhahassan2):

I just need help with part d

OpenStudy (mtalhahassan2):

how u get 7%

OpenStudy (anonymous):

no i got 2013. and part d is 7%

OpenStudy (mtalhahassan2):

do u get 2013

OpenStudy (mtalhahassan2):

but how

OpenStudy (anonymous):

I've taken this assignment. and just remember.

OpenStudy (mtalhahassan2):

just help me with part d

OpenStudy (mtalhahassan2):

please

OpenStudy (mtalhahassan2):

@KamiBug

OpenStudy (anonymous):

sorry :(.......dont know

OpenStudy (mtalhahassan2):

ok np

OpenStudy (danjs):

Population at time t = Initial population*(rate)^time P(t) = A*b^t

OpenStudy (mtalhahassan2):

I only need help with part d

OpenStudy (danjs):

rate of 3 percent per year b = 0.03

OpenStudy (mtalhahassan2):

ok

OpenStudy (danjs):

p(t) = A*(0.03)^t Let t = 0 at 1990 , and t=t is the year after 1990, so in 1990 t=0 and A = 45000

OpenStudy (danjs):

sorry 1.03

OpenStudy (mtalhahassan2):

so t be 1990

OpenStudy (danjs):

p(t) = 45000*(1.03)^t t=years after 1990

OpenStudy (danjs):

in 2007, t will be 2007 - 1990 = 17 t=17 in year 2007 calculate p(17) = 45000*(1.03)^(17)

OpenStudy (mtalhahassan2):

oh ok

OpenStudy (mtalhahassan2):

I will do it know

OpenStudy (danjs):

The population will have doubled when it is 90000 Solve 90000 = 45000(1.03)^t for t, remember t=years after 1990

OpenStudy (danjs):

IF th epopulation took 10 years to double, the growth rate would be? solve 90000 = 45000(b)^10 for the rate b.

OpenStudy (danjs):

good?

OpenStudy (danjs):

each question is just asking you to solve for different variables using the same equation

OpenStudy (danjs):

lol i just went back to read, i like @allibricker12 response. You say you need help, then the first reply is 7%, some help. lol how you get that?

OpenStudy (mtalhahassan2):